# Extracting the singularity of the logarithmic partition function

### v1

@@ Abstract
The transition points of the two-dimensional square lattice $J-Q_{2}$ and $J-Q_{3}$ models have been numerically determined as $J_{c}/Q=0.04502$ and $J_{c}/Q=0.67046$, respectively [46, 47, 63, 62]. The numerical results of the DRTE for these two models are presented in Fig. 2 (d) and (g), where we also show the results for the derivative of logarithmic partition function itself in (e) (f) and (h) (i). The presence of a dominant background term makes it difficult for the derivative of the logarithmic partition function to directly reveal the weakly first-order nature of the phase transition. As shown in Fig. 2 (e,f) for $J-Q_{3}$ and (h,i) for $J-Q_{2}$ model, the curves appear smooth across the transition point, even for large system sizes. On the other hand, the DRTEs of both the $J-Q_{3}$ and $J-Q_{2}$ models exhibit clear discontinuities near the phase transition points. The results, by definition, prove that the phase transitions in both the $J-Q_{2}$ and $J-Q_{3}$ models are (weakly) first-order phase transitions. Notably, compared to $J-Q_{3}$ model, $J-Q_{2}$ model exhibits stronger finite-size effects, making it more difficult to determine its first-order nature prior to our study [62]. Our findings demonstrate that even weakly first-order phase transitions can be clearly detected using DRTE, even in normal sizes.
@@
Conclusion and discussions.— We have proposed that the RTE only contains the singular term of logarithmic partition function, which is highly sensitive around phase transition points. Then we have systematically studied the RTE and its derivative around the QCPs in 2D quantum spin systems. We derived universal scaling forms for the RTE and DRTE, which capture the numerical results and yield the critical exponent $\nu$ through data collapse analysis near (2+1)D QCPs. More importantly, by definition, DRTE can clearly detect weak first-order phase transitions associated with deconfined quantum criticality, a task that is highly challenging for traditional local physical quantities.
@@
Notably, RTE and DRTE is a universally applicable quantity because its definition does not depend on specific models or the precise definition of order parameters. Studying them has the potential to characterize phase transitions that are challenging for traditional local physical quantities, such as topological phase transitions, and to discover new universal critical exponents. Our work introduces a new and powerful paradigm for characterizing quantum phase transitions, including continuous phase transitions, first-order phase transitions, and the particularly challenging weakly first-order phase transitions.
@@ Acknowledgements.
We thank the helpful discussions with Cheng-Xiang Ding, Yan-Cheng Wang. Z.W. thanks the China Postdoctoral Science Foundation under Grants No.2024M752898. The work is supported by the Scientific Research Project (No.WU2024B027) and the Start-up Funding of Westlake University. The authors thank the high-performance computing center of Westlake University and the Beijing PARATERA Tech Co.,Ltd. for providing HPC resources.
@@ Appendix A Method
The $n$th order Rényi entropy is defined as $S^{(n)}=\frac{1}{1-n}\ln\mathrm{Tr}(\rho^{n})$ (or $\frac{1}{1-n}\ln\mathrm{Tr}(\rho_{A}^{n})$), where $\rho=e^{-\beta H}/Z$ and $Z=\mathrm{Tr}e^{-\beta H}$ (with $H$ being the Hamiltonian ). Here, $\rho_{A}=\mathrm{Tr}_{B}\rho$ represents the reduced density matrix of subsystem $A$ coupled with an environment $B$. The Rényi entanglement entropy is defined using $\rho_{A}$, while the Rényi thermal entropy (RTE) is based on $\rho$. In this work, we focus on RTE. From the above definitions, it is evident that the R’enyi entropy can be expressed as the ratio of partition functions at two different temperatures: $S^{(n)}=\frac{1}{1-n}\ln\frac{Z(n\beta)}{Z(\beta)^{n}}$ (for simplicity, we consider $S^{(2)}$ in this work).
@@
where $J_{N}=J^{\prime}$ and $J_{0}=J$, others $J_{i}$ locate between the two in sequence. In realistic simulation, we are able to gain any ratio ${Z(J^{\prime})}/{Z(J)}$ in this way. But $Z(n\beta,J^{\prime})/Z^{n}(\beta,J^{\prime})$ can not be obtained. The antidote comes from some known reference point $Z(n\beta,J)/Z^{n}(\beta,J)$. Considering a zero-temperature system and assuming the degeneracy of the ground state is $de$, we can derive that: $S^{(n)}=\frac{1}{1-n}\ln\frac{Z(n\beta)}{Z(\beta)^{n}}=\ln de$+$c((J-J_{c})L^{(1/\nu)})$, where $c$ represents the critical corrections, $J_{c}$ represents the critical point, and $\nu$ is the critical exponent of the correlation length. Therefore, we can use some states with known degeneracy as reference points, such as product states and discrete symmetry-breaking states.
@@
$\begin{split}\frac{dZ(n\beta))/dJ}{Z(n\beta))}&=-n\beta\bigg{\langle}\frac{dH}{dJ}\bigg{\rangle}_{Z(n\beta))}\end{split}$ (S6)
@@
$\frac{dS^{(n)}}{dJ}=\frac{1}{1-n}\bigg{[}-n\beta\bigg{\langle}\frac{dH}{dJ}\bigg{\rangle}_{Z(n\beta))}+n\beta\bigg{\langle}\frac{dH}{dJ}\bigg{\rangle}_{Z}\bigg{]}$ (S7)
@@
where $n_{J_{1}}$ means the number of $J_{1}$ operators including both diagonal and off-diagonal ones. It’s worth noting that there is no “$2$” anymore in the $\langle...\rangle_{Z(n\beta)}$ term, because $n_{J_{1}}$ here contains two replicas’ operators which has already been doubled actually.

### v2

@@ Abstract
The transition points of the 2D square lattice $J-Q_{2}$ and $J-Q_{3}$ models have been numerically determined as $J_{c}/Q=0.04502$ and $J_{c}/Q=0.67046$, respectively [50, 51, 66, 63]. The DRTE results for both models are shown in Figs. 2 (d) and (g), along with the derivatives of the logarithmic partition function in panels (e,f) and (h,i). Due to the dominant background contribution, the derivative of the logarithmic partition function appears smooth across the transition, even for large system sizes, as seen in Figs. 2(e,f) for $J-Q_{3}$ and (h,i) for $J-Q_{2}$ model. This shows that the derivative of the logarithmic partition function can not be directly reveal the weakly first-order nature. In contrast, the DRTE reveals clear discontinuities at the transition points, providing direct evidence that both transitions are (weakly) first-order. Notably, the $J-Q_{2}$ model exhibits stronger finite-size effects than $J-Q_{3}$, which has made its first-order nature more difficult to resolve in previous studies [63]. Our results demonstrate that DRTE can unambiguously detect even weakly first-order transitions, in relatively small systems (up to $L=56$).
@@
Conclusion and discussions.— We proposed that the RTE isolates the singular contribution of the logarithmic partition function, making it highly sensitive near phase transitions. We systematically studied the RTE and its derivative DRTE around QCPs in 2D quantum spin systems. By deriving universal scaling forms, we successfully captured the numerical data and extracted the critical exponent $\nu$ via data collapse near (2+1)D QCPs. Importantly, by definition, the DRTE provides a clear signature to distinguish weakly first-order phase transitions and deconfined quantum critical points even in small sizes, which is highly challenging to detect using conventional local observables.
@@
Notably, RTE and DRTE are universally applicable, as their definitions do not depend on specific models or order parameters. They also have the potential to characterizing phase transitions that are difficult to probe with conventional local observables, such as topological transitions, and may reveal new universal critical behaviors. Our work establishes a general and effective framework for studying quantum phase transitions across a wide range of classes, including continuous, first-order, and particularly elusive weakly first-order transitions.
@@ Acknowledgements.
We thank the helpful discussions with Cheng-Xiang Ding, Yan-Cheng Wang and Wenan Guo. Z.W. thanks the China Postdoctoral Science Foundation under Grants No.2024M752898. The work is supported by the Scientific Research Project (No.WU2024B027) and the Start-up Funding of Westlake University. The authors thank the high-performance computing center of Westlake University and the Beijing PARATERA Tech Co.,Ltd. for providing HPC resources.
@@ Appendix A Method
The $n$th order Rényi entropy is defined as $S^{(n)}=\frac{1}{1-n}\ln\mathrm{Tr}(\rho^{n})$ (or $\frac{1}{1-n}\ln\mathrm{Tr}(\rho_{A}^{n})$), where $\rho=e^{-\beta H}/Z$ and $Z=\mathrm{Tr}e^{-\beta H}$ (with $H$ being the Hamiltonian ). Here, $\rho_{A}=\mathrm{Tr}_{B}\rho$ represents the reduced density matrix of subsystem $A$ coupled with an environment $B$. The Rényi entanglement entropy is defined using $\rho_{A}$, while the Rényi thermal entropy (RTE) is based on $\rho$. In this work, we focus on RTE. From the above definitions, it is evident that the Rényi entropy can be expressed as the ratio of partition functions at two different temperatures: $S^{(n)}=\frac{1}{1-n}\ln\frac{Z(n\beta)}{Z(\beta)^{n}}$ (for simplicity, we consider $S^{(2)}$ in this work).
@@
where $J_{N}=J^{\prime}$ and $J_{0}=J$, others $J_{i}$ locate between the two in sequence. In realistic simulation, we are able to gain any ratio ${Z(J^{\prime})}/{Z(J)}$ in this way. But $Z(n\beta,J^{\prime})/Z^{n}(\beta,J^{\prime})$ can not be obtained. The antidote comes from some known reference point $Z(n\beta,J)/Z^{n}(\beta,J)$: [Assuming that we have calculated the ratios $Z(n\beta,J^{\prime})/Z(n\beta,J)$ and $Z(\beta,J^{\prime})/Z(\beta,J)$ using the method described above, the ratio $Z(n\beta,J^{\prime})/Z^{n}(\beta,J^{\prime})$ can be obtained if $Z(n\beta,J)/Z^{n}(\beta,J)$ is known. For a given zero-temperature phase, and assuming the ground state degeneracy is $d_{e}$, one can derive that: $S^{(n)}=\frac{1}{1-n}\ln\frac{Z(n\beta)}{Z^{n}(\beta)}=\ln de$. Therefore, we can use states with known degeneracies — such as dimer product states (with degeneracy 1 in this work) and discrete symmetry-breaking states (with degeneracy 2 in this work) — as reference points at a given parameter $J$. In this way, $Z(n\beta,J)/Z^{n}(\beta,J)$ can be determined by solving: $\frac{1}{1-n}\ln\frac{Z(n\beta,J)}{Z^{n}(\beta,J)}=\ln de$.]
One might be concerned about how to deal with a Hamiltonian without a state with known degeneracy. In this way, we can introduce a temperature axis starting from infinite temperature and annealing down to the target temperature. The degeneracy at infinite temperature is known as $2^{N}$ (consider an N-spin system with spin-1/2).
@@
$\begin{split}\frac{dZ(n\beta))/dJ}{Z(n\beta))}&=-n\beta\bigg{\langle}\frac{dH}{dJ}\bigg{\rangle}_{Z(n\beta)}\end{split}$ (S6)
@@
$\frac{dS^{(n)}}{dJ}=\frac{1}{1-n}\bigg{[}-n\beta\bigg{\langle}\frac{dH}{dJ}\bigg{\rangle}_{Z(n\beta)}+n\beta\bigg{\langle}\frac{dH}{dJ}\bigg{\rangle}_{Z}\bigg{]}$ (S7)
@@
where $n_{J_{1}}$ means the number of $J_{1}$ operators including both diagonal and off-diagonal ones. It’s worth noting that there is no “22” anymore in the $\langle...\rangle_{Z(n\beta)}$ term, because $n_{J_{1}}$ here contains two replicas’ operators which has already been doubled actually.
Note that the RTE is computed using the reweight-annealing QMC method mentioned above, while the DRTE is obtained directly by separately evaluating the two expectation values involved in Eq. (S9) or Eq. (S10). This implies that the computational cost of evaluating DRTE is roughly twice that of energy calculations. Typically, we perform on the order of one million Monte Carlo steps when computing the energy, and we have adopted the same number of steps in the present study.
Figure S1: $J$-$Q$ models, in which the antiferromagnetic Heisenberg interaction ($J$-term) acts on all nearest-neighbor bonds. (a) The checkerboard $J$-$Q$ model on a square lattice with periodic boundary condition in both directions. The four-spin interaction ($Q$-term) acts on nearest-neighbor bond pairs that form a plaquette. (b) $J-Q_{2}$ model: the four-spin $Q$ interaction covers the entire lattice. (c) $J-Q_{3}$ model: the six-spin $Q$ interaction covers the entire lattice.
Appendix B JQJ-Q models
In this work, we studied three types of $J$-$Q$ models, namely checkerboard $J$-$Q$, $J$-$Q_{2}$, and $J$-$Q_{3}$ models. The Hamiltonian of the checkerboard $J$-$Q$ model on a square lattice is given by [49]
$H=-J\sum_{\langle ij\rangle}P_{ij}-Q\sum_{[ijkl]}P_{ij}P_{kl},$ (S11)
[…]
$H=-J\sum_{\langle ij\rangle}P_{ij}-Q\sum_{[ijklmn]}P_{ij}P_{kl}P_{mn},$ (S12)
, as illustrated in Fig. S1(c). The DRTE results for three models are shown in Figs. S2, along with the derivatives of the logarithmic partition function in panels (b,c), (e,f) and (h,i).
Figure S2: We set $Q=1$ and use $J$ as the tuning parameter. (a) The derivative of the RTE: DRTE $\partial S^{(2)}/\partial J$. An inset shows the data on the right side of the phase transition point. (b) and (c) show the data for the derivative of the partition function at temperatures $\beta$ and $2\beta$ near the transition of the checkerboard $J$-$Q$ model. (d–f) Data for the $J-Q_{3}$ model. (g–i) Data for the $J-Q_{2}$ model.